### v1

Abstract
This paper explores the use of Grover’s Algorithm in the classical rainbow table, uncovering the potential of integrating quantum computing techniques with conventional cryptographic methods to develop a Quantum Rainbow Table Proof-of-Concept. This leverages on Quantum concepts and algorithms which includes the principle of qubit superposition, entanglement and teleportation, coupled with Grover’s Algorithm to enable a more efficient search through the rainbow table. The paper also details on the hardware constraints and the work around to produce better results in the implementation stages. Through this work we develop a working prototype of quantum rainbow table and demonstrate how quantum computing could significantly improve the speed of cyber tools such as password crackers and thus impact the cyber security landscape.
Keywords:
Rainbow Table Hashing Grover’s Algorithm Quantum.
1 INTRODUCTION
Safeguarding of password data is critical in the cybersecurity landscape. In most systems, passwords are stored as cryptographic hash values and not in plaintext. This is done to prevent attackers from possessing these confidential information, in an event of unauthorized access. However, in an event if the attackers manage to catch hold of this information, attackers would then be able devise methods to reverse the hashes using the precomputed table, known as the Rainbow Table. These tables store the result in a hash function, which enables rapid searching time to decipher encrypted password and break hashes [oechslin2003making]. In addition to that, the use of rainbow tables as a tool in password cracking is efficient and effective for attackers [kumar2013rainbow]. With attacks running effectively and efficiently relying on classical search algorithms, our work explores the potential of using the prowess of quantum computing to be used upon discovery.
In this paper, we introduces a novel approach that leverages quantum computing to enhance both the efficiency and the effectiveness of rainbow table attacks. This research aims to compose Proof-of-Concept prototypes of a Quantum rainbow table, merging and applying quantum mechanics principles, together with Grover’s Algorithm [younes2008strength]. This exhibits the abilities of quantum assisted password decryption techniques, emphasizing on the impact quantum computing has on cryptographic analysis. The aim is to explore potential implications of quantum computing on cybersecurity and the use of password cracking tools available.
Classical rainbow tables typically involve a linear search through the table, hence resulting in a worst-case time complexity of $O(N)$, where N is defined as the number of entries in the table. On the other hand, by leveraging on Grover’s Algorithm, a quantum algorithm designed for searching unsorted tables, the time complexity for the searches can be significantly reduced. With comparison to a classical method [khurana2023implementation], Grover’s algorithm runs with a time complexity of $O(\sqrt{N})$ iterations, which is a quadratic improvement in the time taken in the search process. For the implementation of Grover’s Algorithm, the result will indicate if the target entry exists within the database but the algorithm does not return the target’s position or identity. Hence, in order to determine the exact position and identity of the target entry, additional steps are needed [younes2008strength].
To the best of our knowledge, no practical quantum solutions have been explored for MD5 hash rainbow table search. However, similar quantum solutions includes the use of Grover’s Algorithm being explored for a key search for the Advanced Encryption Standard (AES) [grassl2016applying] and the development of quantum key distribution protocol [bennett1992quantum]. There are many quantum properties that are used to improve classical implementation of cybersecurity solutions including works in opaque quantum predicate used in hybrid quantum-classical system with a high level of accuracy [balachandran2021quantum]. The closest work that we could find is a theoretical work on the Hellman’s time/memory tradeoff on the Rainbow Table variant [dunkelman2024quantum]. These are the motivation for this work to develop a working prototype for quantum rainbow table so that theoretical works in this space has an implementation for validation.
In QIris, we use an hybrid quantum-classical approach, combining Grover’s algorithm together with classical pre-processing checks on buckets for matches. Leveraging on Grover’s Algorithm to narrow down the search space, thereafter, performing an classical check on the bucket to determine the specific match. This hybrid approach exhibits the quantum speedup for search while addressing the position-location limitation using classical computations, ensuring that QIris is able to find and identify the correct plaintext in an quantum-efficient manner.
1.1 Tools Used
QIris is developed using Python with the packages, shown in Table 1.
Table 1: Packages and Libraries Used
No. Packages Version Libraries Used
1 qiskit 1.1.0 quantum_info
transpiler
circuit
2 qiskit_aer 0.14.2 Aer
3 numpy 1.26.4 -
4 tweedledum 1.1.1 -
1.2 Limitations and Assumption
Throughout the development of QIris, we faced substantial hardware and processing constraints, resulting in the implementation of 4 qubits for the computation. To adhere to the constraints of quantum circuits, it is necessitated to employ a simplified alternative of hashing [pearson1990fast], which leads to a higher hash collision. In a perfect situation without the hardware constraints, the number of qubits would have no limitations and there would not be a need to implement any further hashing and the plaintext can be represented in their ASCII binary representation. However, this will require a large number of qubits; for instance a single character would need 8 qubits to be represented in a quantum circuit. To circumvent this limitation keywords were mapped into 4 bit qubit representation using a simplified hash.
In addition to hardware limitations, we worked on Quantum Simulators for this paper instead of real quantum computers. The quantum simulator helps in providing an idealised version of a quantum computer on classical machines to test our work and bypasses the limitations of not having access to quantum hardware. However, this approach would also mean that potential quantum errors and the nuanced realities of physical quantum operations were not accounted for, as the simulators operated under the assumption of ideal quantum behavior. This assumption provided a controlled environment for algorithm development and testing but may not fully represent the challenges encountered on actual quantum hardware.
A limitation of the Grover’s Algorithm observed during implementation is that when there are only two potential outcomes, Grover’s algorithm cannot provide a solution. For such cases, a classical linear search is used in place of Grover’s algorithm.
1.3 Methodology
QIris aims to leverage on the speed advantages of quantum computing to enhance the classical rainbow table approach. As described earlier in section 1, Grover’s Algorithm with a time complexity of $O(\sqrt{N})$, can significantly improve search efficiency. By using Grover’s Algorithm to replace parts of the linear search process, QIris achieves a faster overall time complexity.
The creation of the rainbow tables and the buckets used in QIris is a one-time process and is not factored in the time complexity. Once these tables are established, they remain static. However, the search for a specific hash within these tables is a dynamic process that occurs at runtime, this is factored into the time complexity, where Grover’s algorithm is employed to accelerate the search. The figure below shows the flow of QIris during runtime.
Figure 1: Methodology of QIris.
1.4 Rainbow Table And Buckets Generation
1.4.1 Rainbow Table Generation
The rainbow table utilized in QIris is generated by precomputing and storing both the initial chain plaintext and the final reduction plaintext into a text file format (.txt) file. To increase the table’s efficacy and diversity, four distinct reduction functions were used in its construction. The table was constructed using four different reduction functions to enhance its effectiveness and diversity. Every reduction function converts the hash value into a plaintext of different lengths, which increases the plaintext’s diversity and raises the likelihood of a successful match. The initial chain plaintext will be hashed using MD5 before the first reduction function is implemented. The base62 encoding method is applied to the alphanumeric characters set utilized in the reduction functions. The numerals 0–9, capital letters A-Z and the lower case letters a-z, make up the characters.
The first reduction function is implemented to reduce the MD5 hash to a 6 characters alphanumeric string. Firstly, by taking the first 8 characters of the MD5 hash and converting them into the integer representation value. This integer is then incremented by 2. The characters are generated by taking the modulus value of the integer with the length of the alphanumeric characters set. To ensure diversity and avoid repetition of characters generated, the integer is divided by the length of the alphanumeric characters set. The code for the first reduction function is shown below:
def reduction_function1(md5_hash):
hash_int = int(md5_hash[:8], 16) + 2
plaintext = ""
for _ in range(6):
plaintext += characters[hash_int % char_len]
hash_int //= char_len
return plaintext
The second reduction function implemented is similar to the first, with a notable variation. The MD5 hash is reduced to a 4 characters alphanumeric string and the integer is incremented by 3 instead of 2. The code for the second reduction function is shown below:
def reduction_function2(md5_hash):
hash_int = int(md5_hash[:8], 16) + 3
plaintext = ""
for _ in range(4):
plaintext += characters[hash_int % char_len]
hash_int //= char_len
return plaintext
The third reduction function implemented is similar to the second, with a notable variation. The MD5 hash is reduced to a 5 characters alphanumeric string and the integer is incremented by 4 instead of 3. The code for the third reduction function is shown below:
def reduction_function3(md5_hash):
hash_int = int(md5_hash[:8], 16) + 4
plaintext = ""
for _ in range(5):
plaintext += characters[hash_int % char_len]
hash_int //= char_len
return plaintext
The fourth reduction function implemented is similar to the third, with a notable variation. The MD5 hash is reduced to a 3 characters alphanumeric string and the integer is incremented by 1 instead of 4. The code for the third reduction function is shown below:
def reduction_function4(md5_hash):
hash_int = int(md5_hash[:8], 16) + 1
plaintext = ""
for _ in range(3):
plaintext += characters[hash_int % char_len]
hash_int //= char_len
return plaintext
The varying increment of the integer in the different reduction functions introduce variation and reduce predictability in the reduction process. This allows a healthy generation of different chains ensuring greater diversity and reducing the collision rate.
1.4.2 Bucket Generations
QIris employs a bucket strategy to streamline the search process by organizing hashes into predefined groups also known as “buckets” in this paper. This strategy reduces the search space that Grover’s Algorithm needs to process. By reducing the number of entries in the search space, the strategy improves the efficiency of the time needed to search through the rainbow table. This method optimizes the overall performance by ensuring that Grover’s Algorithm operates on a smaller search space, thereby accelerating the post-search process.
The strategy utilizes the result of the simplified alternative hashing method [[pearson1990fast]] as mentioned in section 1.2, that hashes the final reduction plaintext into a hash value. A list with numbers ranging from 0 to 65534 is generated and shuffled using a random seed. The random seed is used to ensure reproducibility. A value is initially calculated using the length of the input text modulo 65535. For each character in the input text, this value is recalculated using an index from the list, The index is determined by the sum of the current value and the ASCII value of the character modulo operated on 65535. The final value is then used in the bucket strategy. The code for the simplified alternative hashing method is shown below:
random.seed(44)
permutation_16bit = list(range(65535))
random.shuffle(permutation_16bit)
def hash16bit(txt):
h = len(txt) % 65535
for i in txt:
h = permuation_16bit[(h + ord(i)) % 65535]
return h
As described earlier, the bucket strategy first determines the bucket key using the simplified alternative hashing method and taking the value divided by 16. If a bucket key does not already exist, the bucket key is created. If the bucket key exists, the value modulo by 16 will be appended to the bucket. The code for the bucket strategy is shown below:
end_hashed = hash16bit(end)
bucket_key = end_hashed // 16
if bucket_key not in buckets:
buckets[bucket_key] = []
buckets[bucket_key].append(end_hashed % 16)
1.5 Iterative Implementation
QIris implements a MD5 hash checker to ensure that the hash input is in a valid MD5 format. The code for the MD5 hash checker is shown below:
def is_valid_md5(hash_str):
return re.match(r’^[a-f0-9]{32}\$’, hash_str) is not None
QIris employs classical linear search both before and after applying Grover’s Algorithm. Before applying Grover’s Algorithm, a linear search is performed within the buckets to verify that the bucket key exists. If the bucket key does not exist in the bucket, QIris skips the Grover’s search and subsequent linear search, moving directly to the previous chain, thus saving time. After Grover’s search, if the result is found, a linear search is performed to locate for the given hash and rebuild the chain to retrieve the plain text. If the result is not found, QIris will move on to the previous chain. This process continues until QIris has examined all four chains. If the hash is not found in any of the chains, QIris concludes with the hash being deemed not present in the rainbow table. The code for the implementation is shown below:
funcs = get_reduction_function()
for i in range(len(funcs)):
i += 1
funcs_to_use = funcs[-i:]
remaining_funcs = funcs[:-i]
current_hash = search_hash
for j in range(len(funcs_to_use)):
current_text = funcs_to_use[j](current_hash)
if j != len(funcs_to_use) - 1:
current_hash = md5_hash(current_text)
current_map = hash16bit(current_text)
bucket_key = current_map // 16
if bucket_key not in buckets:
continue
good_states = buckets[bucket_key]
lookup = current_map % 16
if len(good_states) <= 2:
if lookup in good_states:
grover_result = True
else:
grover_result = False
Else:
Grover Search Codes
if Grover Search Counts > 512:
grover_result = True
else:
grover_result = False
if grover_result is True:
index = rainbow_table_end_hashed.index(current_map)
if current_text != rainbow_table_end[index]:
continue
current_text = rainbow_table_start[index]
current_hash = md5_hash(current_text)
for k in range(len(remaining_funcs)):
current_text = remaining_funcs[k](current_hash)
current_hash = md5_hash(current_text)
if current_hash == search_hash:
return current_text
return None
1.6 Quantum Implementation
1.6.1 State Preparation
State preparation is an essential process in quantum computing that involves setting up the quantum system to the desired state before performing any quantum operations. When a quantum circuit is initialized, all qubits are set to the state |0⟩. In a typical state preparation step, qubits are changed to be in desired state, which generally is a superposition of states. In QIris, the state preparation involves setting the search space of which Grover’s algorithm is applied.
Assume that in a four-qubit quantum circuit, the potential measurement outcomes range from binary $0$ to $1$, which correspond to integers 0 to 15. If there are only 7 items (1, 2, 8, 9, 10, 12, 15) in the search space (array elements), the probability of measuring the other 8 items will be 0. Therefore, in this case, the state preparation will be configured to ensure that the measurement outcomes are limited to these 7 items only.
To set the state preparation for the above example, an array [0, 1, 1, 0, 0, 0, 0, 0, 1, 1, 1, 0, 1, 0, 0, 1] with the positions 1, 2, 8, 9, 10, 12, 15 are marked to be part of the superposition representation. The Frobenius norm [sd_frobenius] of the array is then calculated, thereafter converting the elements in the array to a probability amplitude value through the division of each element with the calculated Frobenius norm. Therefore, the superposition for the above array can be written as:
$\frac{1}{\sqrt{7}}|1\rangle+\frac{1}{\sqrt{7}}|2\rangle+\frac{1}{\sqrt{7}}|8\rangle+\frac{1}{\sqrt{7}}|9\rangle+\frac{1}{\sqrt{7}}|10\rangle+\frac{1}{\sqrt{7}}|12\rangle+\frac{1}{\sqrt{7}}|15\rangle$
To derive $\frac{1}{\sqrt{7}}$ as the probability amplitude, the Frobenius norm for the array is calculated, which is $\sqrt{7}$. Thus, every element that is probable in the array is divided by the Frobenius norm to create the final array:
$[0,\frac{1}{\sqrt{7}},\frac{1}{\sqrt{7}},0,0,0,0,0,\frac{1}{\sqrt{7}},\frac{1}{\sqrt{7}},\frac{1}{\sqrt{7}},0,\frac{1}{\sqrt{7}},0,0,\frac{1}{\sqrt{7}}]$
This array is then passed to the Statevector function to generate the state preparation circuit. The concept above is used in QIris during the state preparation and the code used to generate the state preparation is shown below:
norm = np.linalg.norm(data)
data = data / norm
state = Statevector(data)
1.6.2 Grover’s Algorithm
Grover’s Algorithm comprises two main components: the oracle and the amplitude amplification. The oracle marks the computational basis state as the state that it is interested in finding, while the amplitude amplification increases the amplitude of the marked state [8622457].
The oracle generate function is implemented to generate the oracle for the Grover’s Algorithm based on the target given hash lookup value.The input will first be converted to binary representation in little endian format before converting it into a boolean expression which will be password into the PhaseOracle function from qiskit.circuit.library, which will generate the oracle circuit.
The oracle generate function is implemented to generate the oracle for the quantum circuit based on the provided target state. The target state refers to the integer to search for. The target state is first converted to a binary representation string. The string is then reversed to convert it from big endian format to little endian format. The binary string will then be converted into a boolean expression string which will be passed into the PhaseOrcale function to generate the oracle circuit. The code for the oracle circuit is shown below:
def oracle_generate(target_state):
binary_string = format(target_state, ’04b’)[::-1]
boolean_expression = ""
for x in range(len(binary_string)):
if binary_string[x] == "0":
boolean_expression += f"~x{x} & "
else:
boolean_expression += f"x{x} & "
boolean_string = boolean_expression[:-3]
return PhaseOracle(boolean_string)
The algorithm then uses an amplification circuit, also known as the diffusion operator, which amplifies the probability of measuring the correct state. The process involves repeatedly applying the oracle and the amplifier a certain number of times, which is determined by the formula:
$\centering\left\lfloor\frac{\pi}{4}\times\sqrt{N}\right\rfloor\@add@centering$ (1)
where $N$ equals the number of elements in the search space
The GroverAlgo function is implemented to generate the complete Grover’s algorithm circuit and run the circuit through the simulator. Firstly, the Grover operation circuit is generated by passing through the oracle and state preparation circuit to the GroverOperator function. This Grover circuit is then appended to the end of the state preparation, with the number of iterations determined based on the formula above for the optimal number of iterations. At the end of the circuit, all qubits are marked as to be measured. The final Grover’s algorithm circuit is then transpiled and sent to the simulator.
def GroverAlgo(init_state, oracle, iter, bits):
grover = GroverOperator(oracle, state_preparation=init_state)
for _ in range(iter):
init_state.append(grover, range(bits))
init_state.measure_all()
backend = Aer.get_backend(’qasm_simulator’)
compiled_circuit = transpile(init_state, backend)
job = backend.run(compiled_circuit)
counts = job.result().get_counts()
return counts
To determine whether the Grover’s Algorithm has successfully found the target state, we evaluate the result based on the measurement counts. Specifically, the state with the highest count must represent more than 50% of the total shots. In QIris, the total number of shots is set to 1024, so the maximum count for a state must exceed 512 for the algorithm to be considered successful. The code for the evaluation is shown below:
max_count = max(grover_result.values())
if max_count > 512:
grover_result = True
else:
grover_result = False
1.7 Future Works
With this Proof-of-Concept Quantum Rainbow Table, QIris, demonstrates a hybrid approach that integrates classical linear search with quantum computing techniques. Moving forward, several avenues for future work are envisioned. Firstly would be the implementation and exploration by using various hash functions which include SHA1, SHA2, to evaluate on the versatility and performance coupled with different hashing algorithms. The main focus of this would include the speed of quantum computer and the hashing process, ensuring that there is an strong emphasis on comparing the time efficiency of classical and quantum hashing. In addition to that, this open the possibility to explore concepts like quantum salt, where the salt value is being placed in a superposition state, this could push for significant improvements in the security and efficiency of hashing processes. Lastly, more studies will be required to investigate the possibility of these innovations and to assess how to optimize both classical and quantum hashing techniques that are available. Moreover, this could pave the way for more advanced cryptographic methods and applications, leveraging the full potential of quantum computing.
1.8 Conclusion
Our work explores the tip of an iceberg on the integration of quantum computing coupled with the classical techniques to enhance rainbow table, leading to the development of QIris, a Quantum Rainbow Table Proof-of-Concept. Upon using Grover’s algorithm, QIris would run with a search time complexity of $O(\sqrt{N})$ instead of the time complexity of a classical rainbow table of $O(N)$, showcasing the power of Quantum Computing. This hybrid approach of QIris combines classical and quantum methods effectively, circumventing the hardware constraints. We hope that our implementation of the quantum rainbow table will provide researchers in this field with a valuable tool to better study quantum hash attacks
In addition, our work highlights the potential of using quantum computing in cybersecurity, covering a wide range of possible projects mentioned under future works. QIris represents a novel way of applying quantum computing to cryptographic analysis, in hopes to open new avenues for more secure systems as quantum hardware progresses.
[…]